# Meta-learning three-factor plasticity rules for structured credit assignment with sparse feedback

### v1

@@ 2 Method
where the expectation $\langle\cdot\rangle_{S}$ is considered over independent sessions $S$. This requires the computation of the sensitivity of the mean weight update wrt to the plasticity parameters $\frac{\partial\mu^{\,(h)}_{ij}}{\partial\theta_{kl}}$ over training. To that end, we propagate the gradients of the within-trial pre-activations $\mathbf{x}^{t}$, $\boldsymbol{\chi}_{k,\ell}^{t}\in\mathcal{R}^{N}$ (state tangent), the pre-activation trace $\bar{\mathbf{x}}^{t}$, $\boldsymbol{\psi}_{k,\ell}^{t}\in\mathcal{R}^{N}$ (trace tangent), and of the eligibility traces of each synaptic pair $ij$, $e^{t}_{ij}$, ${[\mathbf{z}^{t}_{k,\ell}]}_{ij}$ (eligibility tangent), as well as inter-trial sensitivities of weight matrices (weight matrix tangent), $\mathbf{U}_{k,\ell}^{(h)}$ (c.f. Appendix Sec. C).
@@ Appendix A Plasticity gradient with REINFORCE approximation
$\nabla_{\theta}\langle R\rangle$ $=\nabla_{\theta}\int\pi(\boldsymbol{\Delta}\mathbf{W}\mid\theta)\,R\,\text{d}R$
@@
$=\int\pi(\boldsymbol{\Delta}\mathbf{W}\mid\theta)\,\frac{\nabla_{\theta}\pi(\boldsymbol{\Delta}\mathbf{W}\mid\theta)}{\pi(\boldsymbol{\Delta}\mathbf{W}\mid\theta)}\,R\,\text{d}R$
@@ Appendix B Analysis of dynamics for each network
using a damped Newton method. To avoid identifying duplicate fixed points, we initialized from $200$ random initial conditions and discarded solutions within a distance of $\gamma\leq 1e-5$ of one another. This resulted in a set of unique fixed points per input condition.
@@ Appendix C Tangent-propagation through learning
To be able to compute the gradient of the weight updates with respect to the plasticity parameters, we propagate the gradients of the within-trial pre-activations $\mathbf{x}^{t}$, $\boldsymbol{\chi}_{k,\ell}^{t}\in\mathcal{R}^{N}$ (state tangent), the pre-activation trace $\bar{\mathbf{x}}^{t}$, $\boldsymbol{\psi}_{k,\ell}^{t}\in\mathcal{R}^{N}$ (trace tangent), and of the eligibility traces of each synaptic pair $ij$, $e^{t}_{ij}$, ${[\mathbf{z}^{t}_{k,\ell}]}_{ij}$ (eligibility tangent). Thus we define the following within-trial tangents (sensitivities) with respect to the plasticity parameter $\theta_{k,\ell}$
@@ C.1 Validation of weight updates gradients wrt plasticity parameters
The resulting two versions of the gradients are in very close agreement throughout all $500$ trials (Fig. 2).

### v2

@@ 2 Method
where the expectation $\langle\cdot\rangle_{S}$ is considered over independent sessions $S$. This requires the computation of the sensitivity of the mean weight update wrt to the plasticity parameters $\frac{\partial\mu^{\,(h)}_{ij}}{\partial\theta_{kl}}$ over training. To that end, we propagate the gradients of the within-trial pre-activations $\mathbf{x}^{t}$, $\boldsymbol{\chi}_{k,\ell}^{t}\in\mathcal{R}^{N}$ (state tangent), the pre-activation trace $\bar{\mathbf{x}}^{t}$, $\boldsymbol{\psi}_{k,\ell}^{t}\in\mathcal{R}^{N}$ (trace tangent), and of the eligibility traces of each synaptic pair ijij, $e^{t}_{ij}$, ${[\mathbf{z}^{t}_{k,\ell}]}_{ij}$ (eligibility tangent), as well as inter-trial sensitivities of weight matrices (weight matrix tangent), $\mathbf{U}_{k,\ell}^{(h)}$ (c.f. Appendix Sec. C).
@@ Appendix A Plasticity gradient with REINFORCE approximation
$\nabla_{\theta}\langle R\rangle$ $=\nabla_{\theta}\int\pi(\boldsymbol{\Delta}\mathbf{W}\mid\theta)\,R\,\text{d}\boldsymbol{\Delta}\mathbf{W}$
@@
$=\int\pi(\boldsymbol{\Delta}\mathbf{W}\mid\theta)\,\frac{\nabla_{\theta}\pi(\boldsymbol{\Delta}\mathbf{W}\mid\theta)}{\pi(\boldsymbol{\Delta}\mathbf{W}\mid\theta)}\,R\,\text{d}\boldsymbol{\Delta}\mathbf{W}$
@@ Appendix B Analysis of dynamics for each network
using a damped Newton method. To avoid identifying duplicate fixed points, we initialized from 200200 random initial conditions and discarded solutions within a distance of $\gamma\leq 1e-5$ of one another. This resulted in a set of unique fixed points per input condition.
@@ Appendix C Tangent-propagation through learning
To be able to compute the gradient of the weight updates with respect to the plasticity parameters, we propagate the gradients of the within-trial pre-activations $\mathbf{x}^{t}$, $\boldsymbol{\chi}_{k,\ell}^{t}\in\mathcal{R}^{N}$ (state tangent), the pre-activation trace $\bar{\mathbf{x}}^{t}$, $\boldsymbol{\psi}_{k,\ell}^{t}\in\mathcal{R}^{N}$ (trace tangent), and of the eligibility traces of each synaptic pair ijij, $e^{t}_{ij}$, ${[\mathbf{z}^{t}_{k,\ell}]}_{ij}$ (eligibility tangent). Thus we define the following within-trial tangents (sensitivities) with respect to the plasticity parameter $\theta_{k,\ell}$
@@ C.1 Validation of weight updates gradients wrt plasticity parameters
The resulting two versions of the gradients are in very close agreement throughout all 500500 trials (Fig. 2).